OpenStudy (anonymous):

find limit as x approaches pi/2 of x sec^2 x

OpenStudy (ikram002p):

as you know sec ^2 pi/2 does no exist ,

OpenStudy (ikram002p):

so we need to find a trick for how doing it sec x=1/cos x right ?

OpenStudy (anonymous):

yes and cos pi/2 is 0

OpenStudy (ikram002p):

then lim x sec^2 x = lim x / cos ^2 x do you know LH rule ?

OpenStudy (anonymous):

dont think so

OpenStudy (ikram002p):

LH means for g(x)/h(x) limit g(x) / f(x) = limit g'(x) / f'(x)

OpenStudy (ikram002p):

so for what you have lim x / cos ^2 x = limit x' / (cos^2 x) ' = limit 1/ -2 sin x

OpenStudy (ikram002p):

hence :- limit 1/-2 sin x as n goes to pi/2 = -1/2 :) does that make sense to u ?

OpenStudy (anonymous):

i think so yes

OpenStudy (ikram002p):

:D cool then done thats it :P

OpenStudy (anonymous):

tyvm for the help

OpenStudy (ikram002p):

np , ur wlc more info about LH rule http://en.wikipedia.org/wiki/L%27H%C3%B4pital%27s_rule

OpenStudy (anonymous):

i will take a look now

OpenStudy (dumbcow):

@ikram002p, LH rule does not apply here, the limit is not indeterminate .... 0/0 or inf/inf also the derivative for cos^2 is wrong, it should be ---> 1/-2sincos limit of C/0 = infinity http://www.wolframalpha.com/input/?i=lim+x*sec%5E2+x+as+x-%3Epi%2F2

OpenStudy (anonymous):

it can also be solved by converting term containing "cos"into sin and applying "sinx/x" or "x/sinx" =1.